40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jun 12 02:07:42 2012, Jenne, Update, SUS, PRM oplevs left off 
    Reply  Tue Jun 12 16:33:32 2012, steve, Update, SUS, PRM oplevs servo ON for confirmation 
       Reply  Tue Jun 12 17:04:55 2012, steve, Update, SUS, PRM oplevs servo is still bad PRMolvstillosc.png
          Reply  Tue Jul 10 17:47:55 2012, yuta, Update, SUS, PRM oplevs fixed PRMoplevpitOLTF.pngPRMoplevyawOLTF.pngPRMOplevSpectra20120710.png
          Reply  Mon Aug 6 07:59:33 2012, steve, Update, SUS, PRM oplevs servo is still bad PRMeq.png
             Reply  Mon Aug 6 23:05:22 2012, Jenne, Update, SUS, PRM oplevs servo is fine 
             Reply  Mon Oct 15 08:08:33 2012, steve, Update, SUS, PRM damping restored 
Message ID: 6952     Entry time: Tue Jul 10 17:47:55 2012     In reply to: 6805
Author: yuta 
Type: Update 
Category: SUS 
Subject: PRM oplevs fixed 

I centetered PRM oplev, lowered gain and PRM oplev servo is not oscillating any more.
It is OK for more than a softball practice.

C1:SUS-PRM_OLPIT_GAIN = 0.15  (was 0.5)
C1:SUS-PRM_OLYAW_GAIN = -0.3  (was 0.7)

Openloop transfer function:
  Oplev Pitch: gain ~ 12 at 0.69 Hz resonance
  Oplev Yaw: gain ~ 18 at 0.83 Hz resonance
PRMoplevpitOLTF.pngPRMoplevyawOLTF.png

  I adjusted the gain so that oplev damps resonance as much as possible, but not introduce additional noise. I did no calculation, but just measured OSEM spectra (SUSPIT and SUSYAW). Below, you can see the noise reduces at resonance when oplev servo is on, and not increasing at other frequencies. It was introducing noise before. Someone should do more systematic adjustment of oplev servos for all the optics.

PRMOplevSpectra20120710.png
 

ELOG V3.1.3-